# Learning Program Behavioral Models from Synthesized Input-Output Pairs

### v1

@@ 4.5. Results
Speed of Modelizer training data generation. This graph represents the data generation time distribution for different subjects and training set sizes. The x-axis represents the input types, and the y-axis represents the time in minutes. Every training set size is represented by a different color. For every input type, the generator was producing samples with fixed and increasing complexity (partitioned).
We measured the performance of our input generation pipeline in forming a tokenized training set of generated positive samples of data pairs. Our generators can be configured to ignore all PUT runs when PUT fails to process the synthesized input. In this evaluation, our generators we configured to ensure that the number of requested and produced data pairs was equal. Generation of MathML inputs was a special case, where the input generator was executing the PUT twice to first produce from MathML, and then AsciiMath from . The dataset generation time plotted on Figure 5 includes input generation with grammar fuzzer, check for the uniqueness of produced input, input processing with PUT, tokenization of both synthesized input and produced output, and flushing the formatted dataset to the persistent storage. As Figure 5 highlights, for the majority of subject and input generation strategy combination pairs, it required 2-to-4 minutes to synthesize 10,000 samples, 1.5 to 2 hours to produce 250k samples, and more than 6 to 8 hours to produce a dataset with 1,000,000 samples using an Ubuntu Linux workstation equipped with AMD Ryzen Threadripper 3960X 24-Core processor and 256 GB of RAM. The dataset generation time increases almost linearly according to the number of included records. We also see that input generation with an increasing number of non-terminal expansions takes from 0.3 to 2.5 times longer, and is on average 44% slower. Also, we want to mention that by increasing the complexity of synthesized inputs for SQL and MathML we started to get fewer input samples that PUT can successfully process. While this can be seen as a desired effect for stress testing the application, it significantly slows down the dataset generation process if the end goal is to form a dataset containing only positive samples.
1.5 to 2 hours are required to produce a mid-size dataset containing 250,000 samples using a multicore processor.
RQ5: Model Learning Time
@@
Figure 6. Speed of Modelizer learning models from samples with a fixed and increasing complexity (Partitioned). All model learning was done using a single NVIDIA Geforce RTX 4090 GPU.
@@
According to timing records from the model learning log, it takes Modelizer less than half an hour to learn a model from 10,000 synthesized training pairs using NVIDIA RTX 4090 GPU. The similar models will be trained within 7 to 12 hours using a mid-size dataset with 250,000 training samples. Lastly, it takes Modelizer 22 to 49 hours to train a model using a dataset with 1,000,000 records. Figure 6 demonstrates the model learning time distribution across subjects from the evaluation set, training set sizes, and input generation strategies. While an early program mock can be produced already within 30 minutes using a GPU, a significantly more accurate model will be trained within 10 hours.
We also measured from 1% to 59% slowdown, with an average of 11%, while training models containing inputs generated with fixed or increasing complexity. The training time grows almost linearly with the increase of the training set size, given the automatically discovered hyperparameter configurations per subject from the evaluation set.
@@
Training a Modelizer model can be done on a common PC in a few hours.
RQ6: Real-World Data
Our next research question is RQ6: How well do the synthesized samples represent the real-world data?
The quality of synthesized samples depends on the presence and correctness of the corresponding expansion rules that encode data-relevant features in the grammar. We evaluated the quality of the synthesized samples by checking for the inclusion of tokens collected from real-world samples into the vocabularies of learned models. We have collected the following statistics per data type used in our evaluation:
Markdown.:
One top-level element out of 12,783 contained a high number of URL placeholders. Unfortunately, our generators, given the selected non-terminal expansion configuration, have not produced sequences containing more than 51 URL declarations within 1 generation cycle. Thus, the corresponding placeholders were not included in the model’s vocabulary, resulting in encoding these tokens as “UNKNOWN” and the model’s inability to fully correctly predict the output.
HTML.:
Similar to Markdown, in one of the top-level HTML elements we have observed the absence of the same set of URL placeholders in the model’s vocabulary. Additionally, from 9 more top-level elements a few extra HTML tags, like <head> </head> <title> </title>, and meta attributes, like data-cites class, were not seen during the data generation. During the data generation, we filter out attributes that do not affect the semantics of the output data. But, we have not manipulated the test data. Thus the model prediction could be semantically correct but insignificantly differ from the reference sequence.
SQL.:
The test data contained 3 test cases that had been accessing more than 3 tables in a single query. The synthesized queries did not contain such complex samples, which resulted in the model’s inability to produce a fully matching prediction for these test cases.
KQL.:
2 out of 1,000 queries have accessed more than 3 tables in a single request and used 1 additional feature not seen during the data generation. The model cannot predict the output for these queries.
MathML Formulas.:
In 115 cases out of 1,000 formulas, the token sequence contained the unseen operands and operators. This problem originates from the ability of the custom tokenizer to parse and process these elements due to encoding and lack of the corresponding parser rules. The grammar we have defined for the synthesis of MathML Formulas does not cover all possible encodings for operators and contains only a fraction of the special operand space. Due to this fact, the custom tokenizer also fails to parse these special cases. While our tokenizer already abstracts the numeric values and string literals with the corresponding placeholders, we have not implemented the same abstraction mechanism for the special operands. The better encoding and mapping of operators will also help to represent such uncommon inputs without the necessity to store them in the model’s vocabulary. The last issue is the utilization of more than 18 identifiers within a single formula. To synthesize such patterns the data generation configuration has to be changed.
and AsciiMath Formulas.:
Similar to the MathML case we have detected the both larger than synthesized number of enumerated placeholders and the presence of the unseen operands and operators in the test data in 70 and 73 token sequences respectively.
Python and Expressions.:
No unseen tokens were detected.
MathML to AsciiMath behavior mocking and vice-versa, and MathML to  behavior mocking and vice-versa are representative demonstrations of a correlation between the quality of the grammar used for input generation and the accuracy of learned models. The more features are covered by the grammar, the more program behavior gets tested and leaked, and the better model is capable of accurately mocking the program behavior. A fully trained model continuously failing to predict the output acts as a marker of the missing features in the training data, thus in the input grammar. The failure can be detected using a mechanism described in Section 2.4.
At this point, two scenarios are possible:
  • First, the grammar that has been used in the data generation process was an estimate of the actual protocol specification, and it has to be refined according to newly discovered features.
  • Second, the grammar follows the protocol specification, but the program implementation does not. Then the learned model can be used as an oracle to detect the program’s behavior deviation from the protocol specification.
In the case of predicting conversions of MathML Formulas to and AsciiMath representations, we demonstrated the necessity of the population of the grammar with new features and improvements in the tokenization process. Our study of the Pandoc markup converter with Modelizer detected a deviation from the CommonMark Spec Version 0.29 (MacFarlane, 2019). Once we extended our grammar declaration with Extended Markup Syntax, the model accuracy significantly improved. Thus, we think that our models can detect behavior deviations and can assist in the protocol specification refinement.
Altogether, we can conclude that the synthesizing data is suitable to form a representative model vocabulary for the significant part of the real-world data. However, according to the evaluation results that we observed, the absence of feature distribution information encoded as part of grammar or mined from real-world data at the synthesized time cannot guarantee accurate model training. Still, the training with synthesized data allows learning basic patterns and relationships between the input and output tokens, which can be further fine-tuned on the real-world data using just a few representative examples (few-shot training) to achieve the desired accuracy.
The performance of Modelizer depends on the accuracy and completeness of the producing grammar.
RQ7: Usage of Synthetic Data
Our next research question is RQ7: How much synthetic data is required for the behavior mocking of black-box programs? To answer these questions, we need to distribute the training set sizes into three clusters:
Small:
consisting of 5k, 10k, and 50k samples;
Medium:
containing 100k, and 250k samples; and
Large:
with 500k and 1M samples.
The results are summarized in Table 4 and Table 5. We see that the best results for models trained only on synthetic data were achieved with Medium-sized models. They scored the highest number of Exact Matches while also maintaining higher values for Close Matches and BLEU scores. However, after a few-shot fine-tuning with real data, the best-performing models were trained with Large datasets. Such a pattern is representative of half of the subjects we tested. More aggregated statistics for the remaining subjects can be found in Table 3.
Small and medium data sets suffice to obtain good Modelizer models.
RQ8: Generating Synthetic Data
Let us now go to RQ8: How should synthetic data be generated to achieve the highest accuracy of the learned models? To answer this question, we are going to explore the impact of the non-terminal expansion strategy on the accuracy of the learned models and refer to the representative results of Markdown to HTML conversion shown in Table 4 and HTML to Markdown conversion shown in Table 5. As we can see, increasing the complexity of synthesized tokens does not improve the conversion accuracy when the model is trained only on synthetic data. When we fine-tuned the model on a small set number of additional real-world samples that were not part of the test set, we measured an insignificant increase in prediction accuracy.
@@
Without providing the feature distribution by assigning expansion probabilities, we assume the fuzzing budget gets wasted on randomly chosen expansions that do not align with the feature patterns observed in the real-world data. We can conclude that the efficient learning strategy would be to generate the mid-size synthetic dataset with a fixed number of non-terminal expansions, which, according to our results, is sufficient to cover the feature mappings between input and output, and then perform a few-shot fine-tuning on the real-world data, if available. This will achieve the highest or almost the highest possible accuracy of the learned models while maintaining the lowest possible time and resource consumption.
@@
The highest Modelizer performance is obtained by combining a medium-sized synthetic data set with fine-tuning on real-world data.
RQ9: Hyperparameters
@@
Our next research question concerns hyperparameters: How does Modelizer learn the hyperparameters for each data pair combination?
While Modelizer makes no assumptions on the availability of the real-world data at the training time, it still can perform hyperparameter optimization. For that, it needs a fixed set of test cases that are used to evaluate the model’s performance by computing the cross-entropy loss function after every search trial. The optimizer tries to minimize the value of the loss criterion and prunes non-promising trials to further speed up the search process. In our experiments, we reused the same set of test cases for forward and inverse models, which allowed us to study the reusability of discovered hyperparameters for the inverse models. Given the behavior mocking accuracy we have observed during the evaluation of real-world data, we think that running a hyperparameter parameter search on synthesized data is a suitable approach for cases when sufficient quality and quantity of real-world inputs are not available at the training time. According to the results of the hyperparameter search in Table 6, the optimizer managed to find an individual configuration for every translation pair type and the resulting number of model parameters is small, ranging between 2.7 and 10 million parameters for the given subjects.
Hyperparameter optimization is automatic and efficient.
@@
RQ10: Reusability of Hyperparameters
In this research question, we want to check the reusability of the hyperparameters found with forward models for training the inverse models. To answer this question, we need to take a look at the results of Hyperparameter Search presented in Table 6. While the optimal embedding size value equal to 256 and the number of encoder layers equal to 1 were common for all input-output pairs, the rest of the hyperparameters were different. Additionally, we can mention that the best models were mainly trained with the lowest Learning Rate value; Linear Layer Size has to be at least four times larger than the Embedding Size; and, for the same data type pairs, inverse models require mostly two times more decoder layers than forward models. We can conclude that the hyperparameters for the forward models can serve as a lower bound for the hyperparameter search for the inverse models.
Hyperparameters from forward models can be reused for inverse models.
RQ11: Effect of Placeholders
Let us now turn to the effect of design decisions, and start with RQ11: How does the inclusion of additional enumerated placeholders into the model’s input/output vocabularies impact behavior learning?
The inclusion of the additional enumerated tokens into both model’s input/output vocabularies will impact the model’s configuration setup and increase the number of required neurons. Since the training data does not contain these injected tokens, the model will not be able to predict them. Modelizer will deactivate dead neurons, and they will not be used for the prediction. In cases, when the real-world samples contain the injected placeholders, the corresponding neurons still can be revived using the model fine-tuning. Such a technique is a cheap workaround to the problem of the incompatible feature distribution between the training and prediction inputs, when the input/output feature distribution is unknown at data generation time, or the protocol specifications that PUT implements get changed over time affecting only the quantity, but not the variety of these features. It helps to either completely avoid or delay the necessity to retrain the model, which is a time and resource-consuming process.
For several subjects that we tested like bidirectional conversion of the AsciiMath formulas to the representation and vice versa, the MathML formulas to the AsciiMath representation and vice versa, MathML formulas to the representation, the quantity of the synthesized placeholders given the used configuration of non-terminal nodes expansion for data generation was insufficient to represent all symbolic values that can be found in the test set. The preliminary oversizing of the vocabularies with the additional placeholders helps to avoid the model’s failure to predict test cases that contain unseen placeholders. Thus, if the selected hardware used for model learning can tolerate the additional memory allocation caused by the inclusion of additional placeholder tokens into the model’s vocabulary, we recommend using this technique to handle the processing of previously unseen input and output data, which is encoded with additional features, in a combination with fine-tuning.
Placeholders can significantly improve Modelizer performance.
RQ12: Effect of Tokenizers
Let us turn to tokenizers. How should the model input/output be tokenized to achieve the best learning performance? Should the pre-trained tokenizers used instead of domain-specific tokenizers, considering the higher development effort and cost?
The performance of the model is highly dependent on the quality of the generated tokens. The more tokens encode the same input/output data the more complex the model becomes, the more time it requires to produce the output, and the higher the chance it has to fail predicting a correct token sequence. Complex models require more powerful resources for functioning. At the same time, having a very small vocabulary also limits the model’s capability to learn a mapping between input and output correctly. In our work, we suggested using domain-specific tokenizers, which are developed by the domain experts and are capable of tokenizing the input/output data with the matching granularity. The custom tokenization pipeline gives full control over the tokenization process and allows to encoding of the input/output data in the most suitable way. The custom tokenizers are also capable of handling domain-specific data types and formats, which are not currently supported by the pre-trained tokenizers.
@@
Table 7. Comparison of the tokenization quality between implemented domain-specific tokenizers and selection of pre-trained tokenizers.
@@
The pre-trained tokenizers are shipped with the state-of-the-art NLP models BERT and Llama3. Yellow cells indicate an evaluation run with a custom domain-specific tokenizer, blue cells indicate the evaluation runs of the BERT tokenizer, and purple cells indicate the evaluation runs of the Llama3 Tokenizer.
@@
We compared the tokenization quality of the pre-trained tokenizers that are shipped with the state-of-the-art NLP models BERT (Devlin et al., 2019) and Llama3 (AI@Meta, 2024) with the domain-specific tokenizers that were implemented for each subject. We have computed the average length of the tokenized sequences and the vocabulary size per subject. To simplify the comparison we derived additional metrics, the length ratio and vocabulary ratio, which are computed as the ratio of the average length and vocabulary size of the pre-trained tokenizer to the domain-specific tokenizer. The results of our experiments are presented in Table 7. The evaluation was conducted on 1000 real-world samples for each subject, except for Markdown and HTML, where the tokenization was performed on all top-level elements extracted from 100 real-world documents. All inputs that were processed by each tokenizer were augmented by the PlaceholderProcessor module, which replaced the numeric and string literals with the matching placeholders.
@@
While the pre-trained tokenizers managed to formulate a smaller vocabulary for half of the evaluated subjects, the average length of the tokenized sequences was significantly higher than the one produced by the domain-specific tokenizers. We demonstrate an example of the tokenization of the HTML element by the domain-specific and pre-trained tokenizers in Table 8. In this case, the domain-specific tokenizer produces a sequence consisting of 18 tokens, BERT produces 58 tokens, and Llama3 produces 48 tokens. Both BERT and Llama3 split the HTML tags and placeholders into multiple sub-tokens, which increases the length of the tokenized sequence. The domain-specific tokenizer, on the other hand, produces a shorter sequence of tokens that are more meaningful and easier to interpret.
The results of the evaluation in Table 7 show that the domain-specific tokenizers outperform the pre-trained tokenizers in terms of the average length of the tokenized sequences. The pre-trained tokenizers, on the other hand, do not require additional development effort, can tokenize multiple subjects, and are easily interchangeable. We still think that, currently, the domain-specific tokenizers can achieve better tokenization granularity, thus they are more suitable for better-quality mocking. At the same, we think that once a new type of pre-trained tokenizer that focuses on the tokenization of the program-specific data is released, program behavior modeling will become more accessible and easier to implement. Still, we currently suggest using the pre-trained tokenizers as a lower bound for the domain-specific tokenizers.
By repeatedly analyzing the token vocabulary for the test set and using the knowledge of domain experts and feedback from the PUT, one can potentially quicker come up with a domain-specific tokenizer implementation that will outperform the generic pre-trained tokenizers. We are currently working on the automation of this process and plan to reduce the development cost and effort of such tokenizers while keeping the tokenization quality high.

### v2

@@ 4.5. Results
Markdown Sequence Length Frequency Distribution. The x-axis represents the length of different token sequences, and the y-axis denotes the frequency count of each token length. Every data split is represented by a different color.
Figure 6. MathML Sequence Length Frequency Distribution. The x-axis represents the length of different token sequences, and the y-axis denotes the frequency count of each token length. Every data split is represented by a different color.
[…]
MathML Sequence Length Frequency Distribution. The x-axis represents the length of different token sequences, and the y-axis denotes the frequency count of each token length. Every data split is represented by a different color.
To explain such a performance boost we examined the token sequence length frequency for data structures used in our experiments. As expected, the synthesized data differs from real-world data in terms of the number of tokens that these data structures contain. For example, the token sequence length frequency distribution for Markdown that is depicted in Figure 5 shows that the longest recorded synthetic Markdown structure has 368 tokens, with a mean length of 26 tokens. When it comes to real-world data, the longest processed sequence has 598 tokens, with a mean length of 12 tokens. Also, for MathML, whose token sequence length distribution is depicted in Figure 6, we observed a similar pattern. The synthetic data generator produced MathML structures that were tokenized into sequences of up to 159 tokens, while the mean sequence contained 48 tokens. However, real-world MathML structures from the test set contained up to 1038 tokens with a mean MathML formula consisting of 151 tokens. We conclude that synthetic inputs produced by Modelizer generators are shorter than test samples, which affects the model’s ability to generalize to longer real-world inputs. One way to solve this issue is to tweak the input generator’s configuration to produce longer sequences by recursively applying more node expansions from the input grammar. However, a larger expansion budget can be wasted on irrelevant expansions due to random grammar expansion picks at the input generation time. While this process can be guided, for example, by assigning expansion probabilities to rules of interest, it still requires prior knowledge of input feature distributions inferred from at least a few program inputs and incorporation of this knowledge into grammar(Soremekun et al., 2022). The generated dataset will then match the properties of seen inputs, and models trained on such a dataset will still struggle to generalize to dissimilar inputs. Alternatively, fine-tuning pre-trained models on real-world samples could be a more generic solution to the problem. It does not require the regeneration of training data followed by model re-training. Instead, it helps to adapt the capabilities of trained models that contain foundational knowledge about input-output relations to new conditions at a low cost.
Sequence length discrepancy between synthetic and real-world data can be smoothed using few-shot fine-tuning.
For a small number of test samples, fine-tuned models still struggled to accurately convert input to output and vice-versa. The inspection of these cases has shown that they contain unseen features that are not part of the grammar used in input synthesis, for example, features present in the input syntax extensions such as Markdown Extra (Fortin, 2003) or GitHub Flavored Markdown (GitHub Flavored Markdown Specification, 2019). Similarly, all models that had been trained on MathML Formulas used as model input or output (MathML to , to MathML, MathML to AsciiMath, AsciiMath to MathML) have shown the lowest prediction accuracy. When tested on real-world data, these models behaved as outliers because the MathML language allows multiple encodings for the semantically equivalent operators and operands, and the grammar we used for the input generation covers only a fraction of these definitions. This means that selected grammar does not cover all the capabilities of the  PUT, and thus, samples with such features are not present in the training data. If a tokenizer relies on the subject’s syntactical structure learned from grammar, it will also fail to parse these cases.
We also evaluated the quality of the synthesized samples by checking for the inclusion of tokens collected from real-world samples into the vocabularies of learned models. Here are the statistics per data type used in our evaluation:
Markdown.:
One top-level element out of 12,783 contained a high number of URL placeholders. Our generators did not create sequences with over 51 URL declarations per fuzzing round using the chosen non-terminal expansion configuration. Consequently, these placeholders were missing from the model’s vocabulary, leading to their encoding as UNKNOWN tokens and hampering accurate output prediction.
HTML.:
In one top-level HTML element, we observed the absence of matching URL placeholders in the model’s vocabulary. Additionally, in 9 other top-level elements, we noticed the presence of extra HTML tags, like <head>, </head>, <title>, </title>, and meta attributes like the data-cites class, which were not seen during the synthetic data generation because input generator was removing HTML attributes that do not semantically affect data representation. In these cases, the model’s predictions were semantically correct but slightly differed syntactically from the reference sequence.
SQL.:
The test data contained three test cases that had been accessing more than three tables in a single query, whereas the synthesized queries lacked such complexity, limiting the model’s ability to generate matching predictions.
KQL.:
Two out of 1,000 queries have accessed more than three tables in a single request and used one extra feature not seen during the data generation. The model was unable to accurately predict the output for these queries.
MathML Formulas.:
In 115 cases out of 1,000 formulas, the token sequence contained the unseen operands and operators. Our grammar for synthesizing MathML Formulas covers only a fraction of operands and specifies a single encoding per operand type. Thus, the MathML tokenizer also can not parse some operands found in test data due to unseen encoding. An abstraction mechanism for operands, similar to numeric values and string literals, was not implemented. Another issue is the utilization of more than 18 identifiers in a single formula, which was uncovered by the selected input generation configuration.
LaTeX and AsciiMath Formulas.:
We found inconsistencies in placeholder quantities between the train and test sets, similar to MathML, with unseen operands and operators in 70 and 73 test data token sequences, respectively.
Python and Expressions.:
No unseen tokens were detected.
This investigation demonstrates that the quality of learned models depends on the completeness of synthesizing grammar and the similarity of input feature distributions between training and testing sets. The more input features are covered in the grammar, the more program behavior is tested and leaked, and the better the model is at accurately mocking program behavior. A fully trained model that frequently fails to predict the output acts as a marker of the missing features in the training data and, thus, in the input grammar. This means that the grammar that is used in the data generation process is an estimate of the actual protocol specification, and it needs to be refined according to newly discovered features, which could also be mined automatically from the  PUT (see Section 6.10), which also involves the generation of additional samples with new features and the population of the tokenizer vocabulary with new tokens. Since grammar refinement is a significant change to the learning process, it also enforces either the model training restart or modifications to the embedding layer of the trained model to support new tokens and post-training on newly synthesized samples. Otherwise, if the grammar exactly follows the protocol specification it means that the program implementation does not. For example, our case study of the Pandoc markup converter with Modelizer detected deviations from the CommonMark Spec Version 0.29 (MacFarlane, 2014). The model accuracy significantly improved once we extended our grammar declaration with extended syntax support. Thus, we think that our models can assist in detecting behavior deviations from standard and protocol specifications refinement tasks since the learned model acts as a test oracle. Both types of failures can be automatically detected using a mechanism described in Section 2.4.
To further cover the input feature inconsistency problem, we plotted token frequency distribution for evaluated subjects. In contrast to the majority of synthetic data, a few real-world test inputs contained an unusual number of certain placeholder tokens. For example, Figure 7 plot shows us that synthetic input generators do not generate more than 65 TEXT placeholders in a single Markdown structure, with a significant fraction of synthesized structures containing only up to 10 TEXT placeholders. At the same time, our analysis has shown that there are 2 Markdown structures in the real-world test set with 100 TEXT placeholders and almost 400 structures with 25 TEXT placeholders. Also, real-world structures contain from 4.2 to 54.8 times more alignment tokens (depending on the type) than synthetic equivalents, which are instead saturated more with blockquote elements. Even worse disproportion is monitored with MathML (Figure 8) where first synthetic structures contain fewer identifiers and numbers per synthesized formula than in real-world formulas, completely omit some operand types found in real-world formulas, and contain an enormous amount of MathML tag declarations mostly shaped as recursive structures consisting of opening and closing <mn>, <mrow>, <mi> tags. While tweaks to the number of recursive expansions performed by a grammar fuzzer suffice to improve the quality of synthesized Markdown structures and minimize the gap between synthetic and real-world documents, this will have minimal to zero effect in the case of synthesizing MathML formulas. MathML grammar declaration used in our experiments is a significantly more complex structure with multiple mutually-recursive rules. As a result, simply increasing the number of expansion cycles will generate abnormal and unrealistic structures. Moreover, the given grammar does not define some operators and operands frequently seen in the test samples, which means that both grammar and input generator need adjustments to be able to generate higher-quality inputs. These results once more conclude that the quality of learned models depends on the quality and quantity of generated inputs; thus, it is important to have both correct grammar definitions and targeted configurations of input generators.
Modelizer requires the training data and synthesizing grammar to cover the features in the testing data.
Figure 7. Markdown Token Frequency Distribution. The x-axis represents individual tokens, and the y-axis denotes the frequency of each token in the data split. Each split is represented by a different color.
[…]
Markdown Token Frequency Distribution. The x-axis represents individual tokens, and the y-axis denotes the frequency of each token in the data split. Each split is represented by a different color.
Figure 8. MathML Token Frequency Distribution. The x-axis represents individual tokens, and the y-axis denotes the frequency of each token in the data split. Each split is represented by a different color.
[…]
MathML Token Frequency Distribution. The x-axis represents individual tokens, and the y-axis denotes the frequency of each token in the data split. Each split is represented by a different color.
RQ3: How efficiently can Modelizer learn program models?
@@
Figure 9. Speed of Modelizer learning models from samples with a fixed and increasing complexity (Partitioned).
All model learning experiments were done using a single NVIDIA Geforce RTX 4090 GPU.
@@
To answer this question, we need to measure Modelizer’s learning speed on commodity hardware by repeating our learning experiments multiple times on each subject with different training set sizes and fixed subject-specific model configurationsTable 6. According to timing records from the model learning log, Modelizer needs less than half an hour to learn a model from 10,000 synthesized training pairs for 10 epochs using NVIDIA RTX 4090 GPU. The similar models will be trained within 7 to 12 hours using a mid-size dataset with 250,000 training samples. Lastly, Modelizer needs 22 to 49 hours to train a model using a dataset with 1,000,000 records. Figure 9 shows the model learning time distribution across subjects from the evaluation set, training set sizes, and input generation strategies. While an early program mock can be produced already within 30 minutes using a GPU, a significantly more accurate model will be trained within 10 hours. We also measured a slowdown from 1% to 59%, with an average of 11%, while training models containing inputs generated with fixed or increasing complexity. The training time grows almost linearly with the increase of the training set size, given the automatically discovered hyperparameter configurations per each subject from the evaluation set.
@@
Training a Modelizer model takes a few hours on a PC equipped with a commodity GPU.
@@
While looking for an optimal learning strategy, we studied the impact of the non-terminal expansion strategy on the accuracy of the learned models. As we can see in the representative results of Markdown to HTML (Table 4) conversion and HTML to Markdown (Table 5) conversion, continuously increasing the complexity of synthesized tokens does not significantly improve the conversion accuracy neither when the model is trained purely on synthetic data nor after fine-tuning on real-world samples. Without providing the feature distribution as grammar rule expansion probabilities, the larger expansion budget still gets wasted on randomly chosen expansions that do not align with patterns observed in the real-world data. We conclude that the efficient learning strategy is generating the mid-size synthetic dataset with a fixed number of non-terminal expansions, which, according to our results, is sufficient for learning feature mapping between input and output languages. Then, if available, we perform a few-shot fine-tuning on the real-world data. It achieves near-optimal conversion accuracy with low resource use–a pattern we saw in half of the subjects we tested. More statistics for the remaining subjects can be found in Table 3.
@@
Optimal Modelizer performance is achieved with a medium-sized synthetic dataset and real-world fine-tuning.
@@
RQ4: How does Modelizer learn the hyperparameters for each data pair combination?
@@
Our next research question concerns hyperparameters: How does Modelizer learn the hyperparameters for each data pair combination? While Modelizer makes no assumptions on the availability of data at the training time, it still can perform hyperparameter optimization. For that, it needs a fixed set of test cases that can be generated using grammar, which is used to evaluate different model configurations over multiple trials. The optimizer tries to minimize the value of the loss criterion and automatically prunes non-promising configurations to further speed up the search process. In our experiments, we reused the same set of test cases to discover optimal configurations for both forward and inverse models, which allowed us to study the reusability of discovered hyperparameters for the inverse models as well. While the optimal embedding size value and the number of encoder layers were common for all input-output pairs, the rest of the hyperparameters differed. The best models were mainly trained with the smallest Learning Rate value; Linear Layer Size has to be at least four times larger than the Embedding Size; and, for the same data type pairs, inverse models require mostly two times more decoder layers than forward models. Thus, we can conclude that the hyperparameters for the forward models can serve as a lower bound for the hyperparameter search for the inverse models. Given the behavior mocking accuracy we have observed during the evaluation of real-world data, we think that running a hyperparameter search on synthesized data is a suitable approach for cases when sufficient quality and quantity of real-world inputs are not available at the training time. According to the results of the hyperparameter search in Table 6, the optimizer managed to find an individual configuration for every translation pair type, and the resulting number of model parameters is small, ranging between 2.7 and 10 million parameters for the given subjects.
Hyperparameter optimization performed on synthetic data improves the quality of learned models.
RQ5: What is the effect of tokenizer selection on learning performance?
Let us turn to tokenizers. How should the model input/output be tokenized to achieve the best learning performance? Considering the higher development costs, should the pre-trained tokenizers be used instead of custom tokenizers? The model’s performance heavily relies on token quality. Longer token sequences necessitate the use of more complex models, which leads to longer output generation times and a rise in prediction errors. Such models also need more processing power, while models trained from small vocabulary have limited learning flexibility. We propose using custom semantically aware tokenizers developed by domain experts to achieve optimal token sequence length—vocabulary size ratios by leveraging data structure knowledge.
@@
Table 7. Comparison of the tokenization quality between custom-implemented tokenizers and selection of pre-trained tokenizers.
@@
Yellow cells = Custom Tokenizer; Blue cells = BERT Tokenizer; purple cells = Llama3 Tokenizer.
@@
We compared the tokenization quality of the pre-trained tokenizers that are shipped with the state-of-the-art NLP models BERT (Devlin et al., 2019) and Llama3 (AI@Meta, 2024) with the custom structure-aware tokenizers that were implemented for each subject. We have computed the average length of the tokenized sequences and the vocabulary size per subject. To simplify the comparison, we derived additional metrics, such as the length ratio and vocabulary ratio, which are computed as the ratio of the average length and vocabulary size of the pre-trained tokenizer to the structure-aware tokenizer. The results of our experiments are presented in Table 7 and show that the custom structure-aware tokenizers outperform the pre-trained tokenizers in terms of the average length of the tokenized sequences. On the other hand, the pre-trained and unsupervised tokenizers do not require additional development effort, can tokenize multiple subjects, and are easily interchangeable. The evaluation was conducted on 1000 real-world samples for each subject, except for Markdown and HTML, where the tokenization was performed on all top-level elements extracted from 100 real-world documents. All inputs that were processed by each tokenizer were augmented by the PlaceholderProcessor module, which replaced the numeric and string literals with the matching placeholders. We think that custom structure-aware tokenizers can achieve better tokenization granularity and are thus better suited for program behavior mocking tasks.
@@
While the pre-trained tokenizers managed to formulate a smaller vocabulary for half of the evaluated subjects, the average length of the tokenized sequences was significantly higher than the one produced by the structure-aware tokenizers. We demonstrate an example of the tokenization of the HTML element by the structure-aware and pre-trained tokenizers in Table 8. In this case, the structure-aware tokenizer produces a sequence consisting of 18 tokens, BERT produces 58 tokens, and Llama3 produces 48 tokens. Both BERT and Llama3 split the HTML tags and placeholders into multiple sub-tokens, increasing the tokenized sequence length. On the other hand, the structure-aware tokenizer produces a shorter sequence of tokens that are more meaningful and easier to interpret.
It is possible to dynamically infer a structure-aware tokenizer for many program input-output pairs by randomly mutating program inputs and analyzing changes in program execution and program output. We are working on automating this process to reduce the usage effort of such tokenizers while maintaining high tokenization quality.